### v1

@@ 4.3. Success of Multiple COBRAs
Lemma 10.
If the initial network graph is an Erdős–Rényi graph $G(n,p)$ with $p\geq\frac{C_{p}\log n}{n}$, multi-COBRA passes all $T$ phases and each individual COBRA walk covers the whole graph in $O(T\sqrt{\log n})$ rounds with probability at least $1-O(\frac{1}{n})$.
Proof.
By Lemma 6 we know that $G^{\prime}$ the graph we obtain from $G$ after adding self loops - has $1-\lambda_{2}(G^{\prime})\geq\frac{1}{14}$ with probability at least $1-O(\frac{1}{n^{2}})$. Therefore, according to Theorem 1, a COBRA walk succeeds to cover $G^{\prime}$ in $O(\frac{\log n}{(1-\lambda_{2}(G^{\prime}))^{3}})=O(T)=O(\log n)$ rounds with probability at least $1-O(\frac{1}{n^{2}})$. Hence, by the union bound, $\delta(G)\leq n$ COBRAs succeed in $O(\log n)$ phases with probability at least $1-O(\frac{1}{n})$ if run independently. Lemma 9 tells us that walks do not interfere with each other with probability at least $1-O(\frac{\log n}{n^{8}})$, meaning that the statement of the current Lemma holds with probability at least $1-O(\frac{1}{n})-O(\frac{\log n}{n^{8}})=1-O(\frac{1}{n})$. ∎
Remark 11.
The analysis in (cooper2016coalescing, ) is done for simple graphs (i.e. graphs not featuring self-loops). However, the arguments apply verbatim, with symbols reinterpreted to mean the number of outgoing edges of a node instead of the number of neighbors.
4.4. Tree Packing and Broadcast
@@
After multi-COBRA passes all $T$ phases, every edge of the graph belongs to at most $O(T\sqrt{\log n})$ subgraphs.
Proof.
@@
At each phase, an edge is assigned to at most $4\sqrt{\log n}$ subgraphs and there are $T$ phases. ∎
Lemma 13.
@@
After multi-COBRA passes all $T$ phases, each subgraph has a diameter of $O(\log n)$.
Proof.
@@
The multi-COBRA runs for $T=O(\log n)$ phases and in each phase, we add to each subgraph only those nodes that are neighbors of the nodes already included. Consequently, the diameter of the subgraph increases by at most two per phase. ∎
@@
In the rest of this section, we will run protocols on all the subgraphs in parallel. To achieve this despite potential congestion (recall that each edge may belong to multiple subgraphs), we again organize the execution into phases. Each phase spans $4\sqrt{\log n}$ rounds, ensuring that messages sent along any shared edge are distributed across the protocols without conflict. Specifically, if a protocol would send a message along an edge in a particular round when executed independently, all such messages from different subgraphs are scheduled within the same phase. This phased execution allows all protocols to proceed in parallel while respecting the edge capacity. As a result, the combined round complexity of the protocols increases by at most a factor of $4\sqrt{\log n}$ compared to running an individual protocol.
We are now ready to prove Theorem 5.
Proof of Theorem 5.
@@
The algorithm goes as follows. First, let nodes share the information of $n$, $\delta$ and $\Delta$. This can be done in $O(D)$ rounds by constructing a BFS tree. Next, every node $v$ adds $\Delta-deg(v)$ self-loops. Then, parties run multi-COBRA for $T$ rounds that by Lemma 10 result with probability at least $1-O(\frac{1}{n})$ in $\delta$ spanning subgraphs $\{S_{i}\}_{i\in[\delta]}$. By Lemma 13, those have diameter $O(\log n)$. Moreover, by Lemma 12 every edge of the graph belongs to at most $4\sqrt{\log n}$ subgraphs.
@@
Now, we launch BFSs on all subgraphs in parallel to turn them into spanning trees. As discussed earlier in this section, this can be done in $O(4\sqrt{\log n}\cdot\max\limits_{i\in[\delta]}D(S_{i}))=O(\log^{1.5}n)$ rounds. As a result of doing so, the weight of every edge could only have decreased, and the diameter of each subgraph at most doubled. ∎
@@
Having a tree packing with the properties described, we can prove Theorem 2 by adding a final piece.
Proof of Theorem 2.
@@
First, build a tree packing from Theorem 5. Then, $s$ evenly distributes messages among trees, so that each tree receives $\frac{k}{\delta}$ messages. After that, in each tree, nodes downcast corresponding messages. The algorithm for downcasting messages $\{m_{1},\ldots,m_{k}\}$ from the root of a single tree works as follows. In the first round, the root sends the first message ($m_{1}$) to all its immediate children. In the second round, the children forward $m_{1}$ to their respective children (the root’s grandchildren), while the root simultaneously sends the second message ($m_{2}$) to its immediate children. This process continues iteratively: in each subsequent round, the root sends the next message ($m_{i}$) to its children, and all other nodes forward the message they received in the previous round to their respective children. This way, for $k^{\prime}$ messages and a tree of height $h$ it takes $h+k^{\prime}-1$ rounds for every node to discover every message.
@@
Multiplying by a congestion factor of $4\sqrt{\log n}$, we get that the round complexity of broadcasting $k$ messages in $\delta(G)$ spanning trees of height $O(D(G))$ is
$O(4\sqrt{\log n}\cdot(\frac{k}{\delta(G)}+D(G)))=\tilde{O}(\frac{k}{\delta(G)}+D(G))$
Remark 14.
@@
In (ghaffari2015distributed, ; chandra2024fast, ), authors consider a problem where initially messages are spread over the network, that is every node possesses a subset of $M$. This seems like a more general version, however, when using a tree packing approach, these two problems are equivalent. The intuition is, nodes can first agree on the distribution of messages among trees, then upcast the messages to the root in their corresponding trees, and finally perform a downcast as described in our paper, all that in $\tilde{O}(D(G)+\frac{k}{\delta(G)})$. For the full proof, a reader is invited to see the proof of Theorem 1 in (chandra2024fast, ).
5. Sketching the terrain
5.1. NP-Hardness
@@
We prove that the multi-message broadcast problem in CONGEST is NP-hard in the centralized setting. To do that, we reduce the Set splitting problem.
Definition 0 (Set splitting problem).
@@
In our reduction, for simplicity of presentation, we allow edges to have arbitrary bandwidth instead of $\tilde{O}(1)$, since, as we show, this can be simulated in CONGEST. In the reduction the initial set $S$ corresponds to the set $M$ of messages and $s$ has two dedicated children to which it can send $n_{1}$ and $n_{2}$ messages respectively with $n_{1}+n_{2}=k$, simulating the splitting. Deciding how to split messages between these two children is the only “smart“ choice an algorithm should make, all other nodes are just forwarding messages they receive. For the full version of the proof, please see Appendix B.
5.2. Straightforward Lower Bounds are not Enough
Figure 1. An example where diameter and minimum cut are not telling. Here edge labels denote bandwidth.
It is tempting to argue for an approximation factor of an algorithm by comparing its round complexity to two straightforward lower bounds: $D(G)$ and $\frac{k}{minCut(G)}$. Unfortunately, those are not sufficient as there is an instance (see Figure 1) where $D(G)=O(1)$ and $\frac{k}{minCut(G)}=O(1)$, but the optimal answer is $\Omega(\sqrt{k})$. As for NP-hardness, we consider a more general model where edges might have arbitrary bandwidth, but we show that this can be simulated in CONGEST. For details, please see Appendix A.
Appendix A Straightforward Lower Bounds are not Enough
Figure 2. An example of mapping a graph with arbitrary bandwidths to a graph suitable for CONGEST.
@@ Appendix A Straightforward Lower Bounds are not Enough
Proof idea.
The first inequality is straightforward. We prove the second inequality by induction on the length of the path, that is if there is a path in $G$ from $u$ to $v$ of length $l$, then for any nodes $u^{\prime}$ and $v^{\prime}$ in $u$-clique and $v$-clique respectively, there is a path between $u^{\prime}$ and $v^{\prime}$ in $G^{\prime}$ of length $12l+$ 1. ∎
@@
Consider the original graph $G$ and its corresponding CONGEST graph $G^{\prime}$. Together with set $M$ of messages, they define a multi-message broadcast problem in generalized CONGEST and CONGEST respectively. Let $OPT(G)$ and $OPT(G^{\prime})$ denote the optimal round complexities for $G$ and $G^{\prime}$ respectively. Then $OPT(G)\leq OPT(G^{\prime})$.
Proof.
Consider an execution $E^{\prime}$ for $G^{\prime}$ which achieves $OPT$. We claim that we can build an execution $E$ for $G$, such that for every round $r$ of $E^{\prime}$ and for every $v\in V(G)$, after round $r$ in $E$ $v$ knows all the messages that know the nodes of $v$-clique after round $r$ in $E^{\prime}$. To do so, consider round $r$ and some $v\in V(G)$. Let us say that nodes in $v$-clique in $E^{\prime}$ in round $r$ receive messages $M_{1}$ from the $u_{1}$-clique, messages $M_{2}$ from $u_{2}$-clique and so forth for all neighboring cliques. Then, in $E$ $u_{i}$ sends $M_{i}$ to $v$ satisfying the invariant. Note that $u_{i}$ can do this in terms of the bandwidth by construction of the corresponding CONGEST graph, and in terms of knowing $M_{i}$ by the invariant. ∎
Figure 3. An example graph $G$ where diameter and minimum cut are not telling. Here edge labels denote bandwidth.
@@
First, note that $D(G)=4$, hence by Claim 1, $D(G^{\prime})=O(1)$. Second, notice that $minCut(G)=m$ and the minimal maximal bandwidth of an edge adjacent to some node in $V(G)\setminus\{s\}$ is equal to $m$, therefore, by Claim 2, $m-1\leq minCut(G^{\prime})\leq m$. Finally, note that $\Delta(G\setminus s)=3$, hence by Claim 3, $OPT(G^{\prime})\geq OPT(G)$, where $OPT$ is the optimal round complexity. Therefore, it is sufficient to show that $OPT(G)=\Omega(\sqrt{m})$.
@@ Appendix B NP-hardness
Figure 4. Example of transforming a layered graph with arbitrary bandwidths into the graph suitable for CONGEST. Here the number of messages $n$ is $3$.
@@
The proof of the other direction proceeds analogously maintaining the invariant that every node $v\in V(G)$ in $E$ after $r$ rounds knows all the messages that $v_{out}$ knows in $E^{\prime}$ after $22r+$ 2 rounds. ∎
Theorem 2.
@@
Figure 5. Graph $G$ with arbitrary bandwidths for which it is NP-hard to optimally solve multi-message broadcast. Some nodes are depicted in gray since they serve no other purpose but to layer the graph and in reasonable executions should only relay messages.
@@
Appendix C Technical proofs
Proof of Lemma 5.

### v2

@@ 4.3 Success of Multiple COBRAs
Lemma 4.10.
If the initial network graph is an Erdős–Rényi graph $G(n,p)$ with $p\geq\frac{C_{p}\log n}{n}$, all COBRA walk cover the graph in $O(T)$ rounds with probability at least $1-O(\frac{1}{n})$.
Proof 4.11.
By Lemma 4.8 we know that $G^{\prime}$ the graph we obtain from $G$ after adding self-loops - has $1-\lambda_{2}(G^{\prime})\geq\frac{1}{14}$ with probability at least $1-O(\frac{1}{n^{2}})$. Therefore, according to Theorem 3.7, a COBRA walk succeeds to cover $G^{\prime}$ in $O\left(\frac{\log n}{(1-\lambda_{2}(G^{\prime}))^{3}}\right)=O(T)=O(\log n)$ rounds with probability at least $1-O(\frac{1}{n^{2}})$. And Claim 1 tells us that while the COBRA walks are not independent, the marginal distribution of each individual process matches the distribution it would have under independent execution. Hence, we can apply a union bound and conclude that $\delta(G)\leq n$ COBRA walks cover $G^{\prime}$ in $O(\log n)$ phases with probability at least $1-O(\frac{1}{n})$.
We remark that the analysis in [12] is done for simple regular graphs (i.e., regular graphs not featuring self-loops). However, the arguments apply verbatim if self-loops are allowed, with symbols reinterpreted to mean the number of outgoing edges of a node instead of the number of neighbors.
4.4 Tree Packing and Broadcast
@@
After multi-COBRA completes all $T$ phases, every edge of the graph belongs to at most $O(\log n)$ subgraphs.
Proof 4.13.
@@
At each phase, each edge $(u,v)$ is assigned to at most $4$ subgraphs (two via tokens from $u$ to $v$, and two from $v$ to $u$), and there are $T=O(\log n)$ phases.
Lemma 4.14.
@@
After multi-COBRA completes all $T$ phases, each subgraph has a diameter of $O(\log n)$.
Proof 4.15.
@@
The multi-COBRA runs for $T=O(\log n)$ phases, and in each phase, we add to each subgraph only those nodes that are neighbors of the nodes already included. Consequently, the diameter of the subgraph increases by at most two per phase.
@@
In the rest of this section, we will run protocols on all the subgraphs in parallel. To achieve this despite potential congestion (recall that each edge may belong to multiple subgraphs), we again organize the execution into phases. Now, each phase spans $O(\log n)$ rounds, ensuring that messages sent along any shared edge are distributed across the protocols without conflict. Specifically, if a protocol would send a message along an edge in a particular round when executed independently, all such messages from different subgraphs are scheduled within the same phase. This phased execution allows all protocols to proceed in parallel while respecting the edge capacity. As a result, the combined round complexity of the protocols increases by at most a factor of $O(\log n)$ compared to running an individual protocol.
We are now ready to prove Theorem 1.4. See 1.4
Proof 4.16 (Proof of Theorem 1.4).
@@
The algorithm goes as follows. First, let nodes share the information of $n$, $\delta$ and $\Delta$. This can be done in $O(D)$ rounds by constructing a BFS tree. Next, every node $v$ adds $\Delta-\textit{deg}(v)$ self-loops. Then, parties run multi-COBRA for $T$ rounds that by Lemma 4.10 result with probability at least $1-O(\frac{1}{n})$ in $\delta$ spanning subgraphs $\{S_{i}\}_{i\in[\delta]}$. By Lemma 4.14, those have diameter $O(\log n)$. Moreover, by Lemma 4.12, every edge of the graph belongs to at most $O(\log n)$ subgraphs.
@@
Now, we launch BFSs on all subgraphs in parallel to turn them into spanning trees. As discussed earlier in this section, this can be done in $O(\log n\cdot\max\limits_{i\in[\delta]}D(S_{i}))=O(\log^{2}n)$ rounds. As a result of doing so, the weight of every edge could only have decreased, and the diameter of each subgraph at most doubled.
@@
Having a tree packing with the properties described, we can prove Theorem 1.3 by adding a final piece.
See 1.3
Proof 4.17 (Proof of Theorem 1.3).
@@
First, build a tree packing from Theorem 1.4. Then, $s$ evenly distributes messages among the obtained trees, so that each tree receives $\frac{k}{\delta}$ messages. After that, in each tree, nodes downcast corresponding messages. The algorithm for downcasting messages $\{m_{1},\ldots,m_{k}\}$ from the root of a single tree works as follows. In the first round, the root sends the first message ($m_{1}$) to all its immediate children. In the second round, the children forward $m_{1}$ to their respective children (the root’s grandchildren), while the root simultaneously sends the second message ($m_{2}$) to its immediate children. This process continues iteratively: in each subsequent round, the root sends the next message ($m_{i}$) to its children, and all other nodes forward the message they received in the previous round to their respective children. This way, for $k^{\prime}$ messages and a tree of height $h$ it takes $h+k^{\prime}-1$ rounds for every node to discover every message.
@@
Multiplying by a congestion factor of $O(\log n)$, we get that the round complexity of broadcasting $k$ messages in $\delta(G)$ spanning trees of height $O(\log n)$ is
$O\left(\log^{2}n+\log n\cdot\frac{k}{\delta(G)}\right).$
Remark 4.18.
@@
In [23, 8], authors consider a problem where initially messages are spread over the network, that is every node possesses a subset of $M$. This seems like a more general version, however, when using a tree packing approach, these two problems are equivalent. The intuition is, nodes can first agree on the distribution of messages among trees, then upcast the messages to the root in their corresponding trees, and finally perform a downcast as described in our paper, all that in $\tilde{O}(D(G)+\frac{k}{\delta(G)})$. For the full proof, a reader is invited to see the proof of Theorem 1 in [8].
5 Sketching the terrain
In this section, we share insights on the multi-message broadcast problem and why it is difficult to solve optimally. First, in Section 5.1, we sketch the proof of its NP-hardness in the centralized setting. Then, in Section 5.2, we argue that one can not hope to design an algorithm for a general graph that completes in $\tilde{O}(D(G)+\frac{k}{\delta(G)})$ rounds, which implies a need for some fundamentally new techniques.
5.1 NP-Hardness
@@
We prove that determining the optimal number of rounds for the multi-message broadcast problem in CONGEST is NP-hard in the centralized setting. To do that, we reduce the Set splitting problem to the multi-message broadcast.
Definition 5.1 (Set splitting problem).
@@
In our reduction, for simplicity of presentation, we allow edges to have arbitrary bandwidth instead of $\tilde{O}(1)$, since, as we show, this can be simulated in CONGEST. In the reduction, the initial set $S$ corresponds to the set $M$ of messages, and $s$ has two dedicated children to which it can send $n_{1}$ and $n_{2}$ messages respectively with $n_{1}+n_{2}=k$, simulating the splitting. Deciding how to split messages between these two children is the only “smart” choice an algorithm should make; all other nodes are just forwarding messages they receive. For the full version of the proof, please see Appendix C.
5.2 Straightforward Lower Bounds are not Enough
Figure 1: An example where diameter and minimum cut are not telling. Here edge labels denote bandwidth.
It is tempting to argue for an approximation factor of an algorithm by comparing its round complexity to two straightforward lower bounds: $D(G)$ and $\frac{k}{\textit{minCut}(G)}$. Unfortunately, those are not sufficient as there is an instance (see Figure 1) where $D(G)=O(1)$ and $\frac{k}{minCut(G)}=O(1)$, but the optimal answer is $\Omega(\sqrt{k})$. As for NP-hardness, we consider a more general model where edges might have arbitrary bandwidth, but we show that this can be simulated in CONGEST. For details, please see Appendix B.
6 Conclusion
In this work, we explored the multi-message broadcast problem within the CONGEST model. We presented an algorithm that achieves universal optimality, up to polylogarithmic factors, for networks modeled as random graphs and extended our results through a lifting technique to general expander graphs, paying an additional factor of mixing time.
Our study introduces several promising avenues for future investigation. One intriguing direction involves developing distributed processes capable of rapidly covering graphs that also maintain composability, that is, efficiently supporting multiple simultaneous executions under congestion constraints. Another important direction is identifying graph properties beyond traditional metrics such as diameter and minimum cut, which would facilitate establishing tight lower bounds. Subsequently, algorithms matching these lower bounds could be devised, advancing both theoretical understanding and practical efficiency in distributed computing.
Acknowledgments
Authors are grateful to Diana Ghinea and Bernhard Haeupler for helping to improve the paper greatly.
Appendix A Technical definitions
The mixing time of a Markov chain with transition matrix $P$ and stationary distribution $\pi$ can be defined as:
Definition A.1 (Mixing time).
$\tau_{mix}=\min\left\{t:\max_{x}\left\|\frac{P^{t}(x,\cdot)}{\pi}-1\right\|_{2,\pi}\leq n^{-100}\right\},$
where
$\left\|\frac{P^{t}_{x}}{\pi}-1\right\|_{2,\pi}=\sqrt{\sum_{y}\left(\frac{P^{t}(x,y)}{\pi(y)}-1\right)^{2}\pi(y)}.$
In the context of random walks on graphs, the state space corresponds to the edges of the graph. At each step, the walk moves — equiprobably — to one of the edges adjacent to the current edge. One can show that the resulting Markov chain admits a uniform stationary distribution.
Remark A.2.
To ensure that a Markov chain has a stationary distribution, it must be aperiodic. This is commonly enforced by making the walk lazy, i.e., staying in the same state with probability $1/2$.
Definition A.3 (Conductance).
Let $G=(V,E)$ be an undirected graph, and let $\deg(v)$ denote the degree of vertex $v$. For any nonempty subset $S\subset V$, define
$\partial S\;=\;\{\,\{u,v\}\in E:u\in S,\;v\notin S\},\qquad\mathrm{vol}(S)\;=\;\sum_{u\in S}\deg(u).$
The conductance of the cut $(S,S^{c})$ is
$\Phi(S)\;=\;\frac{\lvert\partial S\rvert}{\min\bigl{\{}\mathrm{vol}(S),\,\mathrm{vol}(S^{c})\bigr{\}}}.$
The conductance of the graph $G$ is then
$\Phi(G)\;=\;\min_{\begin{subarray}{c}S\subset V\\ 0<\mathrm{vol}(S)\leq\frac{1}{2}\mathrm{vol}(V)\end{subarray}}\Phi(S).$
Definition A.4 (Chernoff bounds).
Let $X=\sum_{i=1}^{n}X_{i}$ be the sum of independent Bernoulli random variables with $\mathbb{E}[X]=\mu$. Then, the following Chernoff bounds hold:
[…]
Appendix B Straightforward Lower Bounds are not Enough
Figure 2: An example of mapping a graph with arbitrary bandwidths to a graph suitable for CONGEST.
@@ Appendix B Straightforward Lower Bounds are not Enough
Proof B.1 (Proof idea).
The first inequality is straightforward. We prove the second inequality by induction on the length of the path, that is if there is a path in $G$ from $u$ to $v$ of length $l$, then for any nodes $u^{\prime}$ and $v^{\prime}$ in $u$-clique and $v$-clique respectively, there is a path between $u^{\prime}$ and $v^{\prime}$ in $G^{\prime}$ of length $12l+$ 1.
@@
Consider the original graph $G$ and its corresponding CONGEST graph $G^{\prime}$. Together with set $M$ of messages, they define a multi-message broadcast problem in generalized CONGEST and CONGEST respectively. Let $\mathrm{OPT}(G)$ and $\mathrm{OPT}(G^{\prime})$ denote the optimal round complexities for $G$ and $G^{\prime}$ respectively. Then $\mathrm{OPT}(G)\leq\mathrm{OPT}(G^{\prime})$.
Proof B.3.
Consider an execution $E^{\prime}$ for $G^{\prime}$ which achieves $\mathrm{OPT}$. We claim that we can build an execution $E$ for $G$, such that for every round $r$ of $E^{\prime}$ and for every $v\in V(G)$, after round $r$ in $E$ $v$ knows all the messages that know the nodes of $v$-clique after round $r$ in $E^{\prime}$. To do so, consider round $r$ and some $v\in V(G)$. Let us say that nodes in $v$-clique in $E^{\prime}$ in round $r$ receive messages $M_{1}$ from the $u_{1}$-clique, messages $M_{2}$ from $u_{2}$-clique and so forth for all neighboring cliques. Then, in $E$ $u_{i}$ sends $M_{i}$ to $v$ satisfying the invariant. Note that $u_{i}$ can do this in terms of the bandwidth by construction of the corresponding CONGEST graph, and in terms of knowing $M_{i}$ by the invariant.
Figure 3: An example graph $G$ where diameter and minimum cut are not telling. Here edge labels denote bandwidth.
@@
First, note that $D(G)=4$, hence by Claim 2, $D(G^{\prime})=O(1)$. Second, notice that $minCut(G)=m$ and the minimal maximal bandwidth of an edge adjacent to some node in $V(G)\setminus\{s\}$ is equal to $m$, therefore, by Claim 3, $m-1\leq minCut(G^{\prime})\leq m$. Finally, note that $\Delta(G\setminus s)=3$, hence by Claim 4, $\mathrm{OPT}(G^{\prime})\geq\mathrm{OPT}(G)$, where $\mathrm{OPT}$ is the optimal round complexity. Therefore, it is sufficient to show that $\mathrm{OPT}(G)=\Omega(\sqrt{m})$.
@@ Appendix C NP-hardness
Figure 4: Example of transforming a layered graph with arbitrary bandwidths into the graph suitable for CONGEST. Here the number of messages $n$ is $3$.
@@
The proof of the other direction proceeds analogously maintaining the invariant that every node $v\in V(G)$ in $E$ after $r$ rounds knows all the messages that $v_{out}$ knows in $E^{\prime}$ after $22r+$ 2 rounds.
Theorem C.2.
@@
Figure 5: Graph $G$ with arbitrary bandwidths for which it is NP-hard to optimally solve multi-message broadcast. Some nodes are depicted in gray since they serve no other purpose but to layer the graph and in reasonable executions should only relay messages.
@@
Appendix D Technical proofs
Proof D.1 (Proof of Lemma 4.6).